Consider an electromagnet capable of producing a maximum magnetic field of (2.50+1) T. A circular loop of wire with a diameter of (4.50+8) cm is placed between the poles of the electromagnet. What is the induced EMF in the loop, if the electromagnet is turned on to full strength in (1.20+14) milli-seconds? Assume the loop is perpendicular to the magnetic field. Calculate the answer in volts (V) and rounded to three significant figures.
Answer #1

magnetic field B =3.50 T

diameter D =12.50 cm

radius (r) =D/2 =6.25 cm

time taken (t) =15.2 ms

From faradays law

induced emf =N*A*dB/dt

induced emf =(1)*pi*(0.0625 m)^2*(3.50 T)/(0.0152 s)

induced emf =2.825754083 V = 2.830 V